OpenStudy (anonymous):

find two consecutive negative odd integers such that the square of the lesser integer is 40 more than the square of the greater integer

OpenStudy (anonymous):

9 ,11. sounds weird :P

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

i mean-9,-11

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

it amounts to 81 and 121

OpenStudy (anonymous):

i think se said consecutive

OpenStudy (anonymous):

wait im sorry buit i dont understand like how you would do that prblem now..ciykd someone explain that? thanks

OpenStudy (anonymous):

ok for for a general formula try -x^2 =(-x+2)^2 + 40

OpenStudy (anonymous):

but then dont the x's cancel out?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u start with -x^2=(-x+2)^2+40 multiply the square ull get -x^2=x^2-4x+4

OpenStudy (anonymous):

sorry -x^2=x^2-4x+4+40

OpenStudy (anonymous):

oh ok thank you guys!!

OpenStudy (anonymous):

wait where does the 4x come from?